# Quantum Computing Education in Latin America: Experiences and Strategies

### v2

@@ Abstract
Quantum computing is a rapidly advancing field facing a significant shortage of qualified experts, creating an urgent demand for skilled professionals. In Latin America, quantum education remains in its early stages, further exacerbating the regional talent gap due to deficiencies in educational infrastructure and limited financial support. This work presents the efforts of the Quantum Computing Research Group at Universidad Nacional de Colombia to integrate quantum computing into higher education using the European Competence Framework for Quantum Technologies, which offers a standardized approach to defining competency requirements and assessing essential skills in the field. We propose introductory courses aligned with the frameworks guidelines, designed to meet industry standards and reach a broad audience. Additionally, we introduce supplementary resources, including team dynamics and evaluation methodologies, to enhance the educational ecosystem. These initiatives aim to create a sustainable and comprehensive quantum education model across Latin America, to close the regional skills gap, foster inclusivity, and prepare a diverse workforce to contribute innovative solutions on the global stage.
@@ I Introduction
Although the community works to raise awareness and provide basic tools to build expertise, university-level courses are essential for individuals to be qualified and competitive on a global scale. While there are courses and master’s degrees at various universities that cover quantum-related subjects, they are neither common nor widely explored, focused specifically on QC, or accessible to a broader audience. Furthermore, the lack of governmental support results in fewer companies and institutions providing a rigorous environment for research and practical application, making it challenging for students to apply theoretical knowledge in real-world scenarios.
According to the $20232023$ Global Quantum Computing Map [8], the United States, China, and Europe are leading the development of QC, while Japan, South Korea, Singapore, and Australia are also active participants in the global race. Venegas-Gomez [9] presents a global effort map for public funding in QT, an effort that has been enhanced by the establishment of national programs worldwide. However, no Latin American country shows significant investments in QC that would warrant notable mention on the map, nor are they included in the global distribution of companies working in QT per country.
@@ I-A The European Competence Framework for Quantum Technologies
The data underscores not only the urgency of preparing the future workforce, but also the need of tailored educational programs addressing emerging demands [10]. It is crucial to understand the specific qualification required by the industry for various job roles to create targeted courses, specialized tracks, and degree programs at universities which enhance students’ competitiveness for quantum-related positions [11]. To achieve this, the European Competence Framework for Quantum Technologies (CFQT) [12] was developed following a study aimed at collecting and identifying the competences, knowledge, and skills relevant for the emerging quantum workforce and its development [5, 10]. This framework establishes standardized competency requirements for diverse roles in QT, ensuring transparency, scalability, and comparability in defining essential skills, enabling consistent assessment across different courses or qualification [13, 14]. Therefore, one viable solution for Latin American countries is to leverage this framework as a guide to ensure a comprehensive and standardized approach to QC education. Addressing this issue, a research group is dedicated to making QC education open and available to all students in Latin America.
I-B UNAL’S Quantum Computing Research Group
The Quantum Computing Research Group at the Universidad Nacional de Colombia (UNAL) is at the forefront of quantum computing research in Latin America. As one of the largest research groups in the country, with over $7070$ students at various educational levels, they are involved in cutting-edge research projects, publications, conferences, and collaborations with industry, working to integrate QC in the region. The group meets regularly to study key research papers, deliver seminars, and host talks covering a wide range of quantum topics, while maintaining strong connections with researchers and institutions in Europe and the United States. What sets this group apart is its commitment to open science and education. As an inclusive community, it welcomes individuals regardless of academic background and is particularly dedicated to making quantum education accessible across Latin America, where resources in Spanish are scarce or often come with a price barrier, as most educational materials—ranging from formal courses, textbooks, and self-paced online courses, to video series, programming tutorials, workshops, and games— are predominantly restricted to the English language [15], further limiting access for non-English speakers in the region. They consistently offer free courses via Twitch and YouTube, which now have thousands of views and subscribers, making high-quality information available. Their course offerings include “Quantum Computing,” “Artificial Intelligence: Exploring Neural Networks,” and “From Qubits to Qudits.” The group is actively engaging in more national and international collaborations, aiming to build a robust quantum ecosystem in Colombia and Latin America. They are planning to expand their course offerings in the future and intend to use the CFQT to ensure that the education provided aligns with global needs.
@@ II Methodology
We will now present the implementation of the European Competence Framework for Quantum Technologies (version 2.5) to develop two initial quantum computing courses. To begin, we will first explore its structure. The framework consists of three main parts. First, a content map Fig. 1 with $88$ domains and $4242$ subdomains, which provides a graphical overview of the structure and topics of QT.
The framework also contains proficiency levels from A1 to C2 (Tab. I), analogously to the Common European Framework of Reference for Languages (CERF) [16] levels for languages, to tailor education to specific target groups’ needs and to determine job profiles. The depth of knowledge required for each level are described, along with the time it takes to achieve it. These levels apply to three proficiency areas that encompass different competences, each being color-coded according to the domains it covers:
  1. I.
    Quantum Concepts.
  2. II.
    QT Hardware and Software Engineering.
  3. III.
    QT Applications and Strategies.
Thus, the first area (green) is associated with domains $11$ and $22$, the second area (blue) corresponds to domains $33$, $44$, $55$, $66$, and $77$, and the third area (red) is correlated with domain $88$ and the application-oriented subdomains from domains $55$, $66$, and $77$.
@@
Using the proficiency levels and areas, the proficiency triangles (Part $22$) illustrate the general qualifications covered by a person or addressed in a course, as shown in Fig. 2. This figure illustrates a qualification profile created using the Qualification Profile Creation Tool, provided by the CFQT authors, where the individual achieved C1, B2, and C2 proficiency levels in the three respective areas.
@@
Finally, Part $33$ consists of qualification profiles (from P1 to P9) that provide an overview of common personal QT-specific qualifications relevant to industry. The framework shows additional insights on the profiles, including descriptions of the highest levels addressed, a general explanation of the profile, example personas and related recommendations.
II-A Proposed Courses and Implementation Strategy
The designed courses include general information, along with their respective descriptions and objectives. Subsequently, in alignment with the framework’s sections, parameters, and specifications, further aspects are detailed, such as the course content and the target qualification profile that students are expected to achieve. This adaptation involves the following key elements:
@@ II-A Proposed Courses and Implementation Strategy
    A selection of topics and subtopics from the $4242$ available options, carefully chosen to align with the course objectives.
@@
  1. 1.
    Basic Summer/Winter Introductory Course:
@@
    Program:
    1. (a)
      Description: This introductory course provides students with an initial exposure to the current state of QT. It offers a broad perspective on how different quantum technologies interconnect (not solely covering QC), and is suitable for non-STEM degrees and professionals from diverse fields working in quantum-related projects, providing basic knowledge to those without a technical background.
    2. (b)
      Objectives:
      • Provide a foundational understanding of QT, including their applications and the challenges facing the field.
      • Offer a broad perspective on career opportunities and job prospects within the quantum field, as well as the personal impact of working in this sector.
      • Familiarize students with tangible experiences through the use of games, visual activities, or simulations to help grasp key concepts.
      • Promote conceptualization and intuitive understanding of QT, rather than focusing on extensive mathematical competence.
    3. (c)
      Adaptation to the CFQT: For this course, the primary focus aligns with the P1 qualification profile and the Quantum Background domain 1 from the content map, which defines the target profile of a General QT aware person. However, to provide a more comprehensive understanding, the course also incorporates selected subdomains related to Quantum Technology Applications (subdomains 5.6, 6.7, 7.4 and 7.6), offering participants an overview of the broader landscape of QT. Additionally, topics from domain 8 are included to highlight impact and awareness, completing a general overview. This approach ensures that learners gain foundational knowledge while appreciating the potential applications across various fields. The description and qualifications are shown in Table II. Fig. 3 illustrates the proficiency triangle and its levels description, and Fig. 4 shows the chosen focus contents for the course.
@@
  2. 2.
    Undergraduate Elective Course:
@@
    • Duration: Sixteen weeks ($44$ hours per week).
@@
    2. (b)
      Objectives:
      • Offer hands-on experience in programming quantum algorithms, bridging the gap between theory and practice.
      • Facilitate interaction with Ph.D. students and industry professionals to gain insights into real-world challenges and how they are addressed.
      • Foster collaborative learning through Q&A sessions and practical exercises, encouraging active student participation.
    3. (c)
      Adaptation to the CFQT: For this course, the combination of the P3 qualification profile with the Quantum Computing and Simulations domain 5 from the content map leads to the target profile of QC literate person. The description and qualifications are shown in Table III. Fig. 5 illustrates the proficiency triangle and its levels description, and Fig. 6 shows the chosen focus contents.
@@
II-B Additional Proposals
Building on the framework’s guidance for course development, we suggest supplementary elements to augment the courses, ensuring they meet the framework’s standards and enhance the learning experience.
  1. a.
    Work Teams: Analogously to the qualification profiles and example personas given in Part $33$ of the CFQT, examples of real work teams could be integrated into the courses to illustrate applicability and provide students with real-life scenarios of QC applications and understand the tangibility of their future work. A starting point is to use the already existing example personas in the framework, built by combining qualification profiles with a selection from the content map to lead to titles that specify the focus, and later new profiles or professions outside the field can be included. Next, we present a work team example, which does not necessarily reflect a real-life situation but serves for illustrative purposes, to demonstrate how the framework’s concepts could be applied to profile team members.
@@ II-B Additional Proposals
    1. (a)
      Application’s Description: The project focuses on leveraging QC techniques to optimize financial portfolios in real-time. Traditional portfolio optimization methods face limitations in handling large datasets and complex optimization criteria.
    2. (b)
      Objective: Develop and implement quantum algorithms that can optimize financial portfolios more effectively than classical methods, to maximize returns while minimizing risks, considering various financial constraints and market dynamics.
    3. (c)
@@
      • Role: Designs, implements, and optimizes quantum algorithms tailored to financial optimization tasks.
      • Collaboration: Works closely with the Financial Market Analyst and Quantum Computing Specialist to ensure algorithmic feasibility and alignment with financial objectives.
      • Responsibilities: Develops quantum algorithms using appropriate quantum programming tools and software stacks, iteratively refining algorithms based on feedback from the Financial Market Analyst.
@@
      • Role: Provides expertise in QC hardware and software, particularly relevant to financial optimization.
      • Collaboration: Works closely with the Quantum Algorithm Computer Scientist to ensure alignment between algorithmic requirements and quantum hardware capabilities.
      • Responsibilities: Advises on the selection and adaptation of QC techniques to suit the requirements of financial optimization tasks.
@@
      • Role: Analyzes financial market trends, evaluates risk-return profiles, and identifies optimization opportunities.
      • Collaboration: Collaborates with the Quantum Algorithm Computer Scientist and Quantum Computing Specialist to provide insights into financial data and market behavior.
      • Responsibilities: Translates financial objectives and constraints into optimization criteria and assists in defining the performance metrics for quantum algorithms.
@@
      • Role: Develops business strategies for the deployment and commercialization of quantum optimization solutions in the finance sector.
      • Collaboration: Collaborates with the Financial Market Analyst to assess market opportunities and define go-to-market strategies.
      • Responsibilities: Conducts market analysis, identifies potential partners or clients, and develops business plans to drive adoption of quantum optimization solutions.
  2. b.
    Workflow Overview:
    Requirements Analysis: The Financial Market Analyst collaborates with stakeholders to define optimization objectives and constraints. Algorithm Development: The Quantum Algorithm Computer Scientist designs and implements quantum algorithms based on defined optimization criteria. Integration and Testing: The Quantum Computing Specialist ensures seamless integration of algorithms with quantum hardware and conducts performance testing. Market Assessment: The Business Strategist evaluates market opportunities, develops business plans, and identifies potential partners or clients.
  3. c.
    Evaluation: Based on the framework’s proficiency level descriptions for the three areas - Quantum Concepts, QT HW and SW Engineering, and QT Applications and Strategies - which detail the expected knowledge and skills, course evaluation tasks can be developed to assess student learning. By integrating these tasks, clear metrics and assessment tools can be established to measure both the knowledge (K) and skills (S) outlined in the framework, ensuring a thorough assessment of student performance.
@@
I. Quantum Concepts (Target level: A1 - Awareness)
Objective: Evaluate students’ ability to reproduce basic quantum concepts and terminology.
@@ I. Quantum Concepts (Target level: A1 - Awareness)
II. QT Hardware and Software Engineering (Target level: A1 - Awareness)
Objective: Evaluate the basic understanding of the functionalities of quantum HW/SW components, and the ability to reproduce basic functionalities of a QT facet and follow basic instructions.
@@ II. QT Hardware and Software Engineering (Target level: A1 - Awareness)
  • Students follow a series of instructions to simulate a basic quantum circuit using a software simulator (e.g., IBM Quantum Experience).
  • Example: Guided hands-on programming practice - Circuits, application of quantum gates (Hadamard, NOT).
III. QT Applications and Strategies (Target level: A1 - Awareness)
Objective: Evaluate students’ ability to recognize the potential of QT applications and critically engage in public discussions about these technologies.
@@ III. QT Applications and Strategies (Target level: A1 - Awareness)
IV. Overall Evaluation Proposal:
The aim is to ensure a balanced evaluation of both theoretical understanding and practical application, in alignment with the framework’s competencies. Slightly more weight is given to the first area, as the goal is to ensure students have a strong theoretical foundation in quantum concepts. This will provide them with the necessary background to better focus on software, hardware, and applications in subsequent courses. Therefore, in each of the three areas, a greater emphasis has been placed on knowledge assessment rather than skills assessment.
@@ IV. Overall Evaluation Proposal:
II-C Future vision
1. Next Steps in Academia: As a starting point, the first proposed course is under consideration for future testing, potentially to be offered online by UNAL’s research group during a semester break. This initial phase may include updates to their existing QC course available on YouTube. A similar approach could be valuable for other universities looking to implement the framework, as feedback from students and faculty during this phase would be crucial for refining content, delivery methods, and evaluating the effectiveness of integrating quantum education into the actual curriculum in the future, offering online and onsite courses that can be taken with the free elective credits available to students, to implement them across a broader range of careers.
Building on the foundation of introductory and basic QC courses, we aim to develop further undergraduate and graduate specialization courses to gain deeper understanding and practical skills in specific areas. The courses would focus on achieving the B proficiency level and qualification profiles from P4 to P6 depending on students’ profile, interests and experience; as reaching the C level and the P7, P8 and P9 profiles typically requires extensive work experience or advanced research, such as a PhD or post-doc. Furthermore, as QT continue to evolve rapidly, it is imperative to establish continuous professional development programs for educators, providing ongoing training and resources to keep instructors abreast of the latest advancements. In addition, training courses based on the CFQT can be offered to students with more advanced knowledge, enabling them to become educators themselves. This approach not only helps increase the number of trained faculty but also facilitates the dissemination of knowledge to a broader student base, thereby expanding the reach and impact of QC education.

### v3

@@ Abstract
Quantum computing is a rapidly advancing field with the potential to drive scientific, educational, and technological development. However, it faces a significant shortage of qualified experts, creating an urgent demand for skilled professionals. In Latin America, quantum education remains in its early stages, further widening the regional talent and access gap due to limited educational infrastructure and financial constraints. This work presents an initiative to integrate quantum computing into higher education in Latin America through the application of the European Competence Framework for Quantum Technologies, which offers a standardized approach to defining competency requirements and assessing essential skills in the field. We propose introductory courses aligned with the framework’s guidelines, designed to meet industry standards and reach a broad audience. Additionally, we introduce supplementary resources, including team dynamics and evaluation methodologies, to enhance the educational ecosystem. These initiatives aim to create a sustainable and comprehensive quantum education model across Latin America, to close the regional skills gap, foster inclusivity, and prepare a diverse workforce to contribute innovative solutions on the global stage.
@@ I Introduction
Although the community works to raise awareness and provide basic tools to build expertise, university-level courses are essential for individuals to be qualified and competitive on a global scale. While there are courses and master’s degrees at various universities that cover quantum-related subjects, they are neither common nor widely explored, focused specifically on QC, or accessible to a broader audience. One of the main challenges in establishing such courses is the lack of trained faculty members with expertise in quantum computing, making it difficult to integrate the topic into curricula. Furthermore, the lack of governmental support results in fewer companies and institutions providing a rigorous environment for research and practical application, making it challenging for students to apply theoretical knowledge in real-world scenarios.
According to the $2023$ Global Quantum Computing Map [8], the United States, China, and Europe are leading the development of QC, while Japan, South Korea, Singapore, and Australia are also active participants in the global race. Venegas-Gomez [9] presents a global effort map for public funding in QT, an effort that has been enhanced by the establishment of national programs worldwide. However, no Latin American country shows significant investments in QC that would warrant notable mention on the map, nor are they included in the global distribution of companies working in QT per country.
@@ I-A The European Competence Framework for Quantum Technologies
The data underscores not only the urgency of preparing the future workforce, but also the need of tailored educational programs addressing emerging demands [10]. It is crucial to understand the specific qualification required by the industry for various job roles to create targeted courses, specialized tracks, and degree programs at universities which enhance students’ competitiveness for quantum-related positions [11]. To achieve this, the European Competence Framework for Quantum Technologies (CFQT) [12] was developed following a study aimed at collecting and identifying the competences, knowledge, and skills relevant for the emerging quantum workforce and its development [5], [10]. This framework establishes standardized competency requirements for diverse roles in QT, ensuring transparency, scalability, and comparability in defining essential skills, and enabling consistent assessment across different courses or qualification [13], [14]. Therefore, one viable solution for Latin American countries is to leverage this framework as a guide to ensure a comprehensive and standardized approach to QC education.
To address this issue, a Latin American research group is committed to making quantum computing education open and accessible to students across the region. Their work focuses on integrating QC into local educational systems, with a strong emphasis on open science and inclusive learning. This is particularly relevant in Latin America, where educational resources in Spanish are limited or often paywalled, as most materials—such as courses, textbooks, programming tutorials, workshops, and games—are predominantly available in English [15], creating an additional barrier for non-English speakers. The group offers free courses via platforms like Twitch and YouTube and is planning to expand their course offerings in the near future. Their aim is to align these initiatives with the European Competence Framework for Quantum Technologies to ensure global relevance.
This study seeks to explore how the CFQT can be implemented to support the development of quantum computing education in Latin America, and what strategies may enhance its applicability in under-resourced educational contexts.
@@ II Methodology
We will now present the implementation of the European Competence Framework for Quantum Technologies (version 2.5) to develop two initial courses. To begin, we will first explore its structure. The framework consists of three main parts. First, a content map (Fig. 1) with $8$ domains and $42$ subdomains, which provides a graphical overview of the structure and topics of QT.
The framework also contains proficiency levels from A1 to C2 (Tab. I), analogously to the Common European Framework of Reference for Languages (CERF) [16] levels, to tailor education to specific target groups’ needs and to determine job profiles.
@@
The depth of knowledge required for each level are described, along with the time it takes to achieve it. These levels apply to three proficiency areas that encompass different competences, each being color-coded according to the domains it covers:
  1. I.
    Quantum Concepts.
  2. II.
    QT Hardware and Software Engineering.
  3. III.
    QT Applications and Strategies.
Using the proficiency levels and areas, the proficiency triangles (Part $2$) illustrate the general qualifications covered by a person or addressed in a course, effectively outlining its learning objectives. Fig. 2 illustrates an example of a proficiency triangle created using the Qualification Profile Creation Tool, provided by the CFQT authors, where the individual achieved C1, B2, and C2 proficiency levels in the three respective areas.
@@
Finally, Part $3$ consists of qualification profiles (from P1 to P9) that provide an overview of common personal QT-specific qualifications relevant to industry. The framework shows additional insights on the profiles, including descriptions of the highest levels addressed, a general explanation of the profile, example personas and related recommendations.
We now introduce the designed courses, which include general information and descriptions. Subsequently, in alignment with the framework’s sections, parameters, and specifications, further aspects are detailed, such as the course content and the target qualification profile that students are expected to achieve. This adaptation involves the following elements:
@@
    A selection of topics and subtopics from the $42$ available options, carefully chosen to align with the course objectives.
@@
The section “Extending the Framework and Additional Strategies for the Latin American Context” includes both adaptations based on the framework and complementary strategies tailored to the region’s needs. Specifically, building on the framework’s guidance for course development, we suggest supplementary elements —work teams and evaluation methodologies— to augment the courses, ensuring they meet the framework’s standards and enhance the learning experience, while also discussing broader strategies that enhance accessibility and strengthen quantum education in Latin America.
II-A Proposed Courses and Implementation Strategy
II-A1 Basic Summer/Winter Introductory Course
@@
  • Description: This introductory course provides students with an initial exposure to the current state of QT. It offers a broad perspective on how different quantum technologies interconnect (not solely covering QC), and is suitable for non-STEM degrees and professionals from diverse fields working in quantum-related projects, providing basic knowledge to those without a technical background. It aims to familiarize students with tangible experiences through the use of games, visual activities, or simulations to help grasp key concepts, promoting conceptualization and intuitive understanding rather than focusing on extensive mathematical competence.
Adaptation to the CFQT:
For this course, the primary focus aligns with the P1 qualification profile (“QT Aware Person”) and the Quantum Background domain 1 from the content map, which defines the target profile of a General QT aware person. However, to provide a more comprehensive understanding, the course also incorporates selected subdomains related to Quantum Technology Applications (subdomains 5.6, 6.7, 7.4 and 7.6), offering participants an overview of the broader landscape of QT. Additionally, topics from domain 8 are included to highlight impact and awareness, completing a general overview. This approach ensures that learners gain foundational knowledge while appreciating the potential applications across various fields. The description and qualifications are shown in Table II. Fig. 3 illustrates the proficiency triangle and its levels description. Fig. 4 shows the chosen focus contents for the course.
@@
II-A2 Undergraduate Elective Course
@@
  • Duration: Sixteen weeks ($4$ hours per week).
@@
Adaptation to the CFQT:
For this course, the combination of the P3 qualification profile (“QT Literate Person”) with the Quantum Computing and Simulations domain 5 from the content map leads to the target profile of QC literate person. The description and qualifications are shown in Table III. Fig. 5 illustrates the proficiency triangle and its levels description. Fig. 6 shows the chosen focus contents.
@@
II-B Extending the Framework and Additional Strategies for the Latin American Context
II-B1 Work Teams
Analogously to the qualification profiles and example personas given in Part $3$ of the CFQT, examples of real work teams are integrated into the courses to illustrate applicability of students’ future work, providing them with real-life scenarios of quantum applications. This approach also offers a broader perspective on career opportunities within the quantum field, allowing students to explore potential job prospects and the personal impact of working in this sector. In Latin America, where the quantum industry is still emerging, it is beneficial to tailor these examples to regional needs and economic sectors where quantum technologies could have a transformative impact.
A starting point is to use the already existing example personas in the framework, built by combining qualification profiles with a selection from the content map to lead to titles that specify the focus, and later new profiles or professions outside the field can be included. The following work team example does not necessarily reflect a real-life situation, but it serves for illustrative purposes to demonstrate how the framework’s concepts could be adapted to profile team members in sectors relevant to Latin America.
@@ II-B1 Work Teams
  1. I.
    Application’s Description: This project focuses on leveraging quantum computing to analyze financial risks in emerging economies. Unlike stable financial markets, Latin American economies often face high inflation rates and currency fluctuations.
  2. II.
    Objective: Develop and implement quantum algorithms to improve risk assessment models for volatile economies and optimize currency exchange strategies in markets with fluctuating rates, while considering various financial constraints and market dynamics.
  3. III.
@@
      Designs and implements quantum algorithms that integrate macroeconomic indicators into quantum-enhanced predictive models, using appropriate quantum programming tools and software stacks. Works with the Financial Market Analyst and the Quantum Computing Specialist to ensure algorithmic feasibility and alignment with financial objectives.
@@
      Provides expertise in QC hardware and software. Adapts quantum models to available cloud platforms (IBM Quantum, AWS Braket) and advises on the selection of QC techniques to suit the requirements of financial tasks. Works with the Quantum Algorithm Computer Scientist to ensure alignment between algorithmic requirements and quantum hardware capabilities.
@@
      Provides expertise in economic trends, inflation, and currency markets, translating financial objectives and constraints into optimization criteria, and defining the performance metrics for quantum algorithms. Collaborates with the Quantum Algorithm Computer Scientist and the Quantum Computing Specialist to provide insights into financial data and market behavior.
@@
      Develops business strategies for commercialization of quantum solutions in the finance sector. Conducts market analysis, identifies potential partners or clients, and studies the feasibility of integrating quantum models into Latin American financial regulations. Collaborates with the Financial Market Analyst to assess market opportunities and define strategies. Also works with government regulators, banks, and Latin American fintech startups to explore real-world applications.
  4. IV.
    Workflow Overview
    • Data Collection and Economic Analysis: The Financial Market Analyst compiles and processes financial data from Latin American economies, including inflation rates, interest rates, and currency exchange trends. This data is used to define risk assessment parameters and optimization constraints for quantum models.
    • Quantum Model Development: The Quantum Algorithm Computer Scientist designs and implements quantum algorithms tailored to financial risk modeling. The Quantum Computing Specialist assists in selecting the most appropriate quantum algorithms and ensures that the models are computationally feasible.
    • Implementation and Testing on Cloud-Based Quantum Platforms: The Quantum Computing Specialist adapts the models for execution on cloud-based quantum platforms, optimizing for hardware constraints. The models undergo an iterative testing phase, in which the outputs are validated against historical financial data. The Financial Market Analyst assesses the accuracy of the quantum predictions and provides feedback for model refinement.
    • Market Integration and Strategic Deployment: The Business Strategist evaluates the potential for commercial adoption of the quantum models in financial institutions. The final models are integrated into financial decision-making platforms, ensuring that they provide tangible value to financial institutions operating in emerging markets.
II-B2 Evaluation
Building on the framework’s guidance for course development, we suggest an evalaution scheme, ensuring it meets the framework’s standards and enhances the learning experience. Based on the proficiency level descriptions for the three areas - Quantum Concepts, QT HW and SW Engineering, and QT Applications and Strategies - which detail the expected knowledge and skills, course evaluation tasks can be developed to assess student learning. By integrating these tasks, clear metrics and assessment tools can be established to measure both the knowledge (K) and skills (S) outlined in the framework.
@@ II-B2 Evaluation
  1. I.
    Quantum Concepts (Target level: A1 - Awareness)
    Objective:
    • Evaluate students’ ability to reproduce basic quantum concepts and terminology.
@@
  2. II.
    QT Hardware and Software Engineering (Target level: A1 - Awareness)
    Objective:
    • Evaluate the basic understanding of the functionalities of quantum HW/SW components, and the ability to reproduce basic functionalities of a QT facet and follow basic instructions.
@@
    • Students follow a series of instructions to simulate a basic quantum circuit using a software simulator (e.g., Qiskit).
    • Example: Guided hands-on programming practice - Circuits, application of quantum gates (Hadamard, Pauli gates).
  3. III.
    QT Applications and Strategies (Target level: A1 - Awareness)
    Objective:
    • Evaluate students’ ability to recognize the potential of QT applications and critically engage in public discussions about these technologies.
@@
  4. IV.
    Overall Evaluation Proposal: The aim is to ensure a balanced evaluation of both theoretical understanding and practical application, in alignment with the framework’s competencies. Slightly more weight is given to the first area, as the goal is to ensure students have a strong theoretical foundation in quantum concepts. This will provide them with the necessary background to better focus on software, hardware, and applications in subsequent courses. Therefore, in each of the three areas, a greater emphasis has been placed on knowledge assessment rather than skills assessment. This is illustrated in Tab. IV.
@@
II-B3 Cloud-based Infrastructure as a Core Component
Given the limited access to real quantum hardware in Latin America, the courses are designed around cloud-based quantum platforms such as IBM Quantum and Amazon Braket, as well as simulation tools like Qiskit and Cirq. This approach ensures that students not only gain experience with real quantum computing resources but also develop a strong foundational understanding through simulations.
By integrating practical exercises and hands-on programming of quantum algorithms, the courses bridge the gap between theory and practice, allowing students to experiment with quantum circuits before working with actual hardware. Additionally, by familiarizing students with industry-standard tools and workflows, the courses enhance their preparedness for careers, whether in research or industry.
II-B4 PhD and Industry Guest Speakers
The courses intend to integrate sessions with national and international PhD researchers and industry professionals, allowing students to gain insights into real-world challenges and how they are addressed. These sessions will also expose students to global career paths and collaboration opportunities, which is particularly valuable for Latin American students who often have limited direct exposure to the quantum industry. Speakers will discuss not only the personal impact of working in QC and the skills required to thrive in this sector but also share their experiences through interactive discussions, Q&A sessions, and mentorship opportunities.
By fostering a deeper connection between academia, industry, and students, this approach helps bridge the gap between Latin America and the global quantum ecosystem, enhancing career prospects and research opportunities in the region.
II-B5 Regional Collaboration